# How Ensembles of Distilled Policies Improve Generalisation in Reinforcement Learning

### v1

@@ 2.1 Policy distillation
See Appendix D.1 for the proof and exact definitions of all the terms. ∎
@@ 2.3.1 Infinite width limit
See Appendix D.2 for the proof and exact definitions of all the terms. ∎
@@ 4 Generalisation through invariance
$S^{\pi^{*}}_{\mathcal{M|}_{C}}$ =$=\{\psi_{S}(g)s|g\in G,s\in\bar{S}\}$
$S^{\pi^{*}}_{\mathcal{M|}_{C_{train}}}$ =$=\{\psi_{S}(b)s|b\in B,s\in\bar{S}\},\quad B\leq G$
@@ 5.1 Reacher with rotational symmetry
Table 1: Performance of distilled policies in the Illustrative CMDP from Figure 1 for different ensemble sizes $N$ (trained under subgroup $B=C_{4}$) and different subgroups $B\leq SO(2)$ (for $N=1$). Shown are the mean and standard deviation for 20 seeds, and in bold are the best returns including those with overlapping 95% confidence intervals.
@@
Train Performance 1.17 $\pm$ 0.001 1.17 $\pm$ 0.001 1.17 $\pm$ 0.002
Test Performance 0.78 $\pm$ 0.111 0.86 $\pm$ 0.084 0.98 $\pm$ 0.132
@@
Train Performance 1.17 $\pm$ 0.003 1.17 $\pm$ 0.001 1.16 $\pm$ 0.002
Test Performance 0.39 $\pm$ 0.0805 0.78 $\pm$ 0.112 1.11 $\pm$ 0.072
@@ 5.2.1 Policy distillation improves generalisation
Table 3: Performance of an ensemble (of size $N$) of policy distillation or behaviour cloning policies on various datasets compared to the PPO+Explore-GO teacher in the Four Rooms environment. Shown are mean and standard deviation over 20 seeds, and in bold are the best returns including those with overlapping 95% confidence intervals (within the same category).
@@
Lastly, we demonstrate in this section that the same insights also hold for a logarithmic behaviour cloning loss for stochastic policies that is widely used in practice (Foster et al., 2024). At the bottom of Table 3, we show that the an ensemble (of size $N=10$), distilled on the Explore-Go dataset, generalises significantly better than a single behaviour cloning agent on the Teacher dataset. Note that the Mixed dataset does considerably worse for behaviour cloning, because the cloned behaviour policy is a mixture of the PPO policy and a pure exploration policy (that achieves zero return).
@@ A.1 Policy distillation & behaviour cloning
$l(\theta,\mathcal{D},\pi_{\beta})=\frac{1}{n}\sum_{s\in\mathcal{D}}||\pi_{\theta}(s)-\pi_{\beta}(s)||_{2}^{2}$ (6)
@@ A.3 The infinite width limit
$m_{t}(x)$ $=\Theta(x,x_{i})[\Theta^{-1}T_{t}]_{ij}y_{i}$
[…]
where we use the Einstein notation convention to indicate implicit sums over the dataset indices $i,j$, h.c. indicates the Hermitian conjugate of the preceding term, $T_{t}=(\mathbb{I}-\exp(\eta\Theta t))$, $\mathcal{K}(x,x^{\prime})=\mathbb{E}_{\theta\sim\mu}[f_{\theta}(x)f_{\theta}(x^{\prime})]$ is the neural network Gaussian process (NNGP) kernel, and $\Sigma_{t}^{(1)}$ and $\Sigma_{t}^{(2)}$ are defined as follows:
@@
For an infinite ensemble $\bar{f}_{t}$, the function equals the mean $m_{t}$ of the Gaussian process: $\bar{f}_{t}(x)=m_{t}(x)$. Note that for $t\to\infty$, the output of the infinite ensemble $\bar{f}_{\infty}$ on the training inputs $\mathcal{X}$ converges to the targets $\mathcal{Y}$:
@@ B.1 Invariance of an ensemble
For completeness we repeat the proof in our own notation in D.2. ∎
@@
For completeness we repeat the proof in our own notation in D.3. ∎
@@ B.2 Performance during testing
For completeness we repeat the proof in our notation in Appendix D.1. ∎
@@ C.1.2 Figure 1 & Table 1
The teacher policy is a handcrafted policy ($a=(-2,2)$ for 12 steps and $a=(2,2)$ afterwards) that is optimal for the starting pose considered in this setting. The neural network consists of three fully connected hidden layers of size $[64,64,32]$ with ReLU activation functions. The policy is distilled with the MSE loss in (6) (which is the same loss as (1) but for vector-valued actions instead of scalar). The exact hyperparameters can be found in Table 4.
@@ C.2 Four Rooms
In the Four Rooms environment (Figure 4), an agent (red triangle) starts in a random location and facing a random direction, and has to move to the goal location (green square) whilst navigating the doorways connecting the four rooms. We modify the original Minigrid implementation a little bit, by reducing the action space from the default seven (turn left, turn right, move forward, pick up an object, drop an object, toggle/activate an object, end episode) to only the first three (turn left, turn right, move forward). Moreover, we use a reward function that gives a reward of 1 when the goal is reached and zero elsewhere, which differs from slightly from the default one that gives 10.9(step countmax $1-0.9*(\frac{\text{step count}}{\text{max steps}})$ for reaching the goal. Lastly, our implementation of the Four Rooms environment allows for more control over the context dimensions, allowing for the construction of distinct training and testing sets. Our version of the Four Room environment can be found at <redacted for review>.
@@ C.2.1 Implementation details
Appendix D Repeated proofs
D.1 Theorem 3 from Maran et al. [2023]
@@ Appendix D Repeated proofs
$\mathcal{W}(T(\cdot|s,a),T(\cdot|\hat{s},\hat{a}))$ $\leq L_{T}(d_{S}(s,\hat{s})+d_{A}(a,\hat{a})),$ $\forall(s,a),(\hat{s},\hat{a})\in S\times A$
$|R(s,a)-R(\hat{s},\hat{a})|$ $\leq L_{R}(d_{S}(s,\hat{s})+d_{A}(a,\hat{a})),$ $\forall(s,a),(\hat{s},\hat{a})\in S\times A.$
@@
Now, let $L_{s}=||Q^{\pi_{2}}(s,\cdot)||_{L}$ be the Lipschitz semi-norm of $Q^{\pi_{2}}(s,a)$ w.r.t. $a$ and define $g_{s}(a)=Q^{\pi_{2}}(s,a)/L_{s}$ with the property $\|g_{s}\|_{L}=1$. This yields
@@
D.2 Lemma 6.2 from Gerken and Kessel [2024]
@@
$\leq||s-\bar{s}||\sum_{l=1}^{L}\mathbb{E}_{\theta\sim\mu}\bigg{[}\bigg{|}\bigg{(}\frac{\partial\pi_{\theta}(s^{\prime})}{\partial\theta^{(l)}}\bigg{)}^{\top}\cdot L(\theta^{(l)})\bigg{|}\bigg{]}$
$=||s-\bar{s}||\hat{C}(s^{\prime})$
@@
$|\bar{\pi}_{t}(s)-\bar{\pi}_{t}\big{(}\psi_{S}(g)s\big{)}|$ $\leq\hat{C}(s)\sqrt{\sum_{i}||s_{\mathtt{p}_{b}^{-1}(i)}-\psi_{S}(g)s||^{2}}\sqrt{\sum_{i}(\sum_{j,k}\Theta_{ij}^{-1}(\mathbb{I}-\exp(-\eta\Theta t))_{jk}y_{k})^{2}}$
$\leq\kappa\hat{C}(s)\sqrt{\sum_{i}||s_{i}||^{2}}\sqrt{\sum_{i}(\sum_{j,k}\Theta_{ij}^{-1}(\mathbb{I}-\exp(-\eta\Theta t))_{jk}y_{k})^{2}}=\kappa C_{\Theta}(s)$
D.3 Lemma B.4 from Gerken and Kessel [2024]

### v2

@@ 2.1 Policy distillation
See Appendix E.1 for the proof and exact definitions of all the terms. ∎
@@ 2.3.1 Infinite width limit
See Appendix E.2 for the proof and exact definitions of all the terms. ∎
@@ 4 Generalisation through invariance
$S^{\pi^{*}}_{\mathcal{M|}_{C}}$ $=\{\psi_{S}(g)s|g\in G,s\in\bar{S}\}$
$S^{\pi^{*}}_{\mathcal{M|}_{C_{train}}}$ $=\{\psi_{S}(b)s|b\in B,s\in\bar{S}\},\quad B\leq G$
@@ 5.1 Reacher with rotational symmetry
Table 1: Performance of distilled policies in the Illustrative CMDP from Figure 1 for different ensemble sizes NN (trained under subgroup $B=C_{4}$) and different subgroups $B\leq SO(2)$ (for $N=1$). Shown are the mean and standard deviation for 20 seeds, and in bold are the best returns including those with overlapping 95% confidence intervals.
@@
Train Performance 1.17 $\pm$ 0.004 1.17 $\pm$ 0.004 1.17 $\pm$ 0.003
Test Performance 0.75 $\pm$ 0.147 0.89 $\pm$ 0.107 1.05 $\pm$ 0.117
@@
Train Performance 1.17 $\pm$ 0.003 1.17 $\pm$ 0.004 1.16 $\pm$ 0.002
Test Performance 0.39 $\pm$ 0.0805 0.75 $\pm$ 0.147 1.11 $\pm$ 0.072
@@ 5.2.1 Policy distillation improves generalisation
Table 3: Performance of an ensemble (of size NN) of policy distillation or behaviour cloning policies on various datasets compared to the PPO+Explore-GO teacher in the Four Rooms environment. Shown are mean and standard deviation over 20 seeds, and in bold are the best returns including those with overlapping 95% confidence intervals (within the same category).
@@
Lastly, we demonstrate in this section that the same insights also hold for a logarithmic behaviour cloning loss for stochastic policies that is widely used in practice (Foster et al., 2024). At the bottom of Table 3, we show that the an ensemble (of size $N=10$), distilled on the Explore-Go dataset, generalises significantly better than a single behaviour cloning agent on the Teacher dataset. Note that behaviour cloning achieves lower performances than distillation, and that BC performs considerably worse on the Mixed dataset. In our definition of behaviour cloning, the student policy learns to imitate whatever policy collected the dataset, by only observing the actions that were actually sampled during collection. Therefore, the BC agent performs worse than the distillation agent, since the latter has access to more information (all the action probabilities of the teacher). On the Mixed dataset, the BC agent clones the behaviour policy that consists of a 50/50 mix of the (optimal) Teacher policy and (suboptimal) pure exploration policy. The resulting cloned behaviour performs even worse than the BC agent trained on the Teacher dataset. In contrast, the policy distillation agent on the Mixed dataset regresses on the action probabilities of the Teacher, on the states encountered by the 50/50 mixture of policies, and therefore has a much better learning target.
6 Discussion and limitations
The experiments in the Four Rooms environment in Section 5.2.1 serve to empirically demonstrate how our insights can be leveraged to significantly enhance the generalisation performance of a reinforcement learning agent through policy distillation. A clear example of this is seen in our ensemble $N=10$, distilled on the Explore-Go dataset, which achieves substantially higher test performance than the original PPO+Explore-Go teacher policy (see Table 3). The potential of policy distillation after training as a tool to improve generalisation was initially identified in Lyle et al. (2022), but we believe the results of this paper provide a more compelling argument and empirical evidence for this phenomenon.
Whether the benefits of performing data augmentation with respect to some symmetry group actually stem from induced invariance or reduced overfitting and other forms of regularisation, is still an ongoing topic of discussion in the literature (Lyle et al., 2020; Shen et al., 2022). To add to this discussion, in Appendix D.1 we measure the invariance of our trained models on the ’Reacher with rotational symmetry’ experiments from Table 1 and plot it against the ensemble size $N$ and subgroup $B\leq SO(2)$. We find that in this particular experiment, the distilled policies do become more invariant as ensemble and subgroup size increase, just as our theory predicts.
Obtaining tight generalisation bounds for neural networks is notoriously challenging (Jiang et al., 2020; Gastpar et al., 2024). Moreover, some of the assumptions for Theorem 1, such as infinitely wide neural networks, are hard to meet in practice. Therefore, we believe the true strength of our theory lies in its ability to identify crucial properties of the dataset distribution and distilled ensemble that are capable of improving generalisation performance. Nonetheless, in Appendix D.2, we analyse how well our results fit the $a+\frac{b}{\sqrt{N}}$ relation identified by our theory. We find that our results reasonably agree with the shape of the theoretical upper bound, suggesting that our bound is not completely vacuous.
Finally, all ensemble members are trained independently, and during inference, can also be evaluated independently (an independent forward pass with an average over the output of the ensemble afterwards). This inherent independence means that both the training and inference processes of the ensemble are parallelisable. If parallelisation is not feasible, the runtime for both training and inference would increase linearly with the ensemble size. It is important to note that all ensemble members are distilled on the same dataset. This means the small number of additional environment steps required to sample this dataset is independent of ensemble size.
@@ 7 Conclusion
Acknowledgments and Disclosure of Funding
We thank Caroline Horsch, Laurens Engwegen and Oussama Azizi for fruitful discussions and feedback. The project has received funding from the EU Horizon 2020 programme under grant number 964505 (Epistemic AI) and was also partially funded by the Dutch Research Council (NWO) project Reliable Out-of-Distribution Generalization in Deep Reinforcement Learning with project number OCENW.M.21.234. The computational resources for empirical work were provided by the Delft AI Cluster (2024) (DAIC) and the Delft High Performance Computing Centre (2024) (DHPC).
@@ A.1 Policy distillation & behaviour cloning
[…]
@@ A.3 The infinite width limit
$m_{t}(x)$ $=\Theta(x,x_{i})[\Theta^{-1}T_{t}]_{ij}y_{j}$
$\Sigma_{t}(x,x^{\prime})$ $=\mathcal{K}(x,x^{\prime})+\Sigma_{t}^{(1)}(x,x^{\prime})-(\Sigma_{t}^{(2)}(x,x^{\prime})+\text{ h.c.})$
where we use the Einstein notation convention to indicate implicit sums over the dataset indices $i,j$, h.c. indicates the Hermitian conjugate of the preceding term, $T_{t}=(\mathbb{I}-\exp(-\eta\Theta t))$, $\mathcal{K}(x,x^{\prime})=\mathbb{E}_{\theta\sim\mu}[f_{\theta}(x)f_{\theta}(x^{\prime})]$ is the neural network Gaussian process (NNGP) kernel, and $\Sigma_{t}^{(1)}$ and $\Sigma_{t}^{(2)}$ are defined as follows:
@@
An infinite ensemble $\bar{f}_{t}$ equals the mean $m_{t}$ of the Gaussian process: $\bar{f}_{t}(x)=m_{t}(x)$. Note that for $t\to\infty$, the output of the infinite ensemble $\bar{f}_{\infty}$ on the training inputs $\mathcal{X}$ converges to the targets $\mathcal{Y}$:
@@ B.1 Invariance of an ensemble
For completeness we repeat the proof in our own notation in E.2. ∎
@@
For completeness we repeat the proof in our own notation in E.3. ∎
@@ B.2 Performance during testing
For completeness we repeat the proof in our notation in Appendix E.1. ∎
@@ Appendix C Experimental details
The code for all the experiments in the main text can be found at https://github.com/MWeltevrede/distillation-after-training.
@@ C.1.2 Figure 1 & Table 1
The teacher policy is a handcrafted policy (a=(2,2)a=(-2,2) for 12 steps and $a=(2,2)$ afterwards) that is optimal for the starting pose considered in this setting. The neural network consists of three fully connected hidden layers of size $[64,64,32]$ with ReLU activation functions. The policy is distilled with the MSE loss in (6) (which is the same loss as (1) but for vector-valued actions instead of scalar). The exact hyperparameters can be found in Table 4.
@@ C.2 Four Rooms
[…]
@@ C.2.1 Implementation details
Appendix D Additional experiments
D.1 Measure of invariance
In order to identify invariance as a key factor for the increased performance in the ’Reacher with rotational symmetry’ environment from Table 1, we will measure how invariant the policy becomes when increasing ensemble or subgroup size. As a measure of invariance we use the variance of the network’s output across the group orbit [Kvinge et al., 2022] (for a completely invariant network, this should be zero). In practical terms, we evaluate each policy on all (360) integer degree rotations of the starting state (the orbit) and compute the total variation (trace of the covariance matrix) over the produced outputs. In Figure 5, we see that for both increased ensemble and subgroup size, as the test performance increases, so does the measure of invariance decrease.
Figure 5: Test return (left axis) compared with the total variation (trace of the covariance matrix) over orbits of the $SO(2)$ group of rotations (right axis) for (a) different ensemble sizes and (b) subgroups $B\leq SO(2)$. The total variation is a measure of how invariant the agent has become with respect to rotations, zero total variation would correspond to perfect invariance. Shown are the mean and 95% confidence intervals over 20 seeds.
D.2 $\frac{1}{\sqrt{N}}$ generalisation bound
In this section, we investigate how well our results fit the $a+\frac{b}{\sqrt{N}}$ form of the generalisation bound from Theory 1. We evaluate different ensemble sizes on the ’Reacher with rotational symmetry’ environment from Figure 1, trained on subgroup $B=C_{4}$ in Table 8. This table includes the results from Table 1, as well as additional results for ensemble sizes $N=1000$ and $N=10.000$. In Table 9, we repeat the results for different ensemble sizes distilled on the Explore-Go dataset in the Four Rooms environment from Table 3, with the addition of ensemble size $N=100$.
In Figure 6, we compare the difference to optimal performance $J^{\pi^{*}}-J^{\hat{\pi}_{\infty}}$ with ensemble size $N$, and plot the best fit to the $a+\frac{b}{\sqrt{N}}$ relation as predicted by our theory. The exact optimal performance $J^{\pi^{*}}$ is not necessarily known, but we estimate it by taking the average train performance instead (which seems to have converged on both environments). The best $a+\frac{b}{\sqrt{N}}$ fit is obtained by computing the optimal linear fit between $y=J^{\pi^{*}}-J^{\hat{\pi}_{\infty}}$ and $x^{\prime}=\frac{1}{\sqrt{N}}$ (and then transforming the solution back to $x=N$).
The results in Figure 6 seem to follow the $a+\frac{b}{\sqrt{N}}$ upper bound to some extend. The upper bound is likely not very tight, but the results seem to indicate the bound is also not vacuous.
(a) Reacher with rotational symmetry
(b) Four Rooms
Figure 6: Difference to optimal performance as a function of ensemble size $N$ in the testing contexts for the (a) Reacher with rotational symmetry and (b) Four Rooms environments. Shown are the mean and 95% confidence intervals over 20 seeds from (a) Table 8 and (b) Table 9, together with the best possible fit for the relation $a+\frac{b}{\sqrt{N}}$ as predicted by our theory.
Table 8: Performance of distilled policies in the Illustrative CMDP from Figure 1 for different ensemble sizes NN (trained under subgroup $B=C_{4}$). Shown are the mean and standard deviation for 20 seeds, and in bold are the best returns including those with overlapping 95% confidence intervals.
Ensemble Size $N$: N=1 N=10 N=100 N=1000 N=10.000
Train Performance 1.17 $\pm$ 0.004 1.17 $\pm$ 0.004 1.17 $\pm$ 0.003 1.17 $\pm$ 0.003 1.17 $\pm$ 0.002
Test Performance 0.75 $\pm$ 0.147 0.89 $\pm$ 0.107 1.05 $\pm$ 0.117 1.13 $\pm$ 0.072 1.17 $\pm$ 0.038
Table 9: Performance of an ensemble (of size NN) of policy distillation or behaviour cloning policies on the Explore-Go dataset in the Four Rooms environment. Shown are mean and standard deviation over 20 seeds, and in bold are the best returns including those with overlapping 95% confidence intervals.
Ensemble Size $N$: N=1 N=10 N=100
Train Performance 0.92 $\pm$ 0.020 0.92 $\pm$ 0.019 0.92 $\pm$ 0.021
Test Performance 0.78 $\pm$ 0.041 0.88 $\pm$ 0.036 0.89 $\pm$ 0.033
Appendix E Repeated proofs
E.1 Theorem 3 from Maran et al. [2023]
@@ Appendix E Repeated proofs
$)\displaystyle\mathcal{W}(T(\cdot|s,a),T(\cdot|\hat{s},\hat{a}$)) $),\displaystyle\leq L_{T}(d_{S}(s,\hat{s})+d_{A}(a,\hat{a}$)), $\forall(s,a),(\hat{s},\hat{a})\in S\times A$
[…]
@@
[…]
@@
E.2 Lemma 6.2 from Gerken and Kessel [2024]
@@
[…]
@@
[…]
E.3 Lemma B.4 from Gerken and Kessel [2024]